OpenStudy (anonymous):

Three added to 4 times a number is equal to 5 less than 6 times the number.

HanAkoSolo (jamierox4ev3r):

all you have to do here is literally translate the question: 3+4n=5-6n

OpenStudy (anonymous):

If you want it in order it would be 4n + 3 = 6n -5

OpenStudy (zzr0ck3r):

im with nightowl 3+4n=6n-5